# Diffraction of large-number whispering gallery mode by boundary straightening with jump of curvature

### v1

@@ 5.1.1 Integral $I_{1}^{--}$
[…]
@@
$(I_{1}^{--})_{0}=\frac{e^{-i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\Bigg{[}\frac{i}{\varepsilon}-\frac{1}{4\mathfrak{t}\,\varepsilon^{2}}+\left(\frac{1}{\sigma}-\frac{1}{\sqrt{\mathfrak{t}}}\right)\frac{1}{2\varepsilon^{3}}\left(1+\delta\right)\Bigg{]}.$ (64)
@@
and $\varepsilon$ is introduced in (52). The need to keep the first three terms of expansion (64) will be clarified later, when adding the contributions of the end point $p=0$ to the integrals $I_{1}^{\pm\pm}$.
@@
Consider merging of the point $p_{1}$ with the end point, which occurs when $\sigma>\sqrt{\mathfrak{t}}$ and the value of $\varepsilon=\sqrt{\mathfrak{t}}-\nu/2\sigma$ is small, see (52). We expand the phase in a small neighbourhood of $p=0$ up to quadratic terms:
@@
Omitting terms of the third order, extending integration from a small neighbourhood of $p=0$ to the half-line $[0,\infty)$ and passing to the variable
@@
see (52). The value $\zeta_{>}^{2}$ equals, up to the constant summand $-2\mathfrak{t}^{3/2}/3$, difference between phase (55) of the wave $u_{2}$ after the rays $\ell_{2}$ pass the caustic and phase (51) of the cylindrical wave. The Fresnel integral in (71) has an imaginary argument, which is atypical of diffraction problems.
@@
The value $\zeta_{<}^{2}$ is equal, up to the constant summand $-2\mathfrak{t}^{3/2}/3$, to difference between the phase of the wave $u_{2}$ before the rays $\ell_{2}$ pass the caustic and the phase of the cylindrical wave, see (51) and (55).
@@
holds true. We expand the phase $f^{--}(p)$ in a small neighbourhood of $p_{0}$ up to cubic terms:
@@
Omitting terms of the fourth order, extending integration from a small neighbourhood of $p_{0}$ to the whole axis and passing to the integration variable $q=(p-p_{0})/2\sigma$, we obtain that the contribution of the critical points of phase to the integral $I_{1}^{--}$ is expressed in terms of the Airy function $v$:
[…]
@@
As above, we expand the phase $f^{--}(p)$ in a neighbourhood of $p_{0}=t-\sigma^{2}$ (which is small now) up to cubic terms, see (79). We move terms of the fourth order to the amplitude, extend the integration to the half-line $[0,\infty)$ and pass to the variable $q=(p-p_{0})/2\sigma$. Keeping in mind that $\sigma-\sqrt{\mathfrak{t}}$ is small, we come up with
$I_{1}^{--}=e^{-i\left(\frac{2\sigma^{3}}{3}+\sigma(\mathfrak{t}-\nu)\right)}\mathcal{I}\left(\widetilde{\nu},\sigma-\sqrt{\mathfrak{t}}\right)\left(1+O\left(\frac{(\sigma-\sqrt{\mathfrak{t}})^{2}}{\sqrt{\mathfrak{t}}}\right)+O\left(\frac{\widetilde{\nu}^{2}+1}{\sqrt{\mathfrak{t}}}\right)\right).$ (82)
@@ 5.1.2 Integrals $I^{++}$, $I^{+-}$ $I^{-+}$
$(I_{1}^{++})_{0}=i\frac{e^{i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\Bigg{[}\frac{i}{\varepsilon}+\frac{1}{4\mathfrak{t}\,\varepsilon^{2}}+\left(\frac{1}{\sigma}+\frac{1}{\sqrt{\mathfrak{t}}}\right)\frac{1}{2\varepsilon^{3}}\left(1+\delta\right)\Bigg{]},$ (83)
@@
Comparing the phase of exponential in (84) with expression (49), we conclude that formula (84) describes the wave $u_{1}$, corresponding to the ray family $\ell_{1}$, after reflection from the straight part of the boundary $S_{+}$. Geometrical spreading of these rays is proportional to the square root of the distance from the point $M^{\prime}$ symmetric to the observation point $M$ with respect to $S_{+}$, to the caustic $C$, see (37). The approximation (83), (84) is applicable in the area where conditions (66) and
@@
cf. (73) and (75). The quantity $\zeta_{*}^{2}$ is equal, up to the constant summand $32\mathfrak{t}^{3/2}/$ 3, to the difference between phase (54) of the wave $u_{1}$ after reflection from the straight part of the boundary and phase (51) of the cylindrical wave. It is easy to check that $\sqrt{(x+a\gamma)/a\gamma}=J(M)/J(O)$, where $J(M)$ and $J(O)$ are the values of geometrical spreading of the rays $\ell_{1}$ at the points $M=(x,\gamma x)$ and $O=(0,0)$, respectively, cf. (76).
@@
$(I_{1}^{+-})_{0}=-\frac{e^{i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}\left(\sqrt{\mathfrak{t}}+\frac{\nu}{2\sigma}\right)}\left(1+O\left(\frac{1}{\mathfrak{t}^{\frac{3}{2}}}\right)\right).$ (87)
Expression (87) has no singularities in area (15).
@@
3. The critical point $p_{4}$ of phase of the integral $I^{-+}$ is negative and lies far from the end point $p=0$ for all values of $\sigma$ and $\nu$. Thus, the asymptotics of $I^{-+}$ is given by the end point $p=0$ contribution. It differs from (87) only by the sign before $\mathfrak{t}^{3/2}$ in exponent and by the constant factor $-i$:
$(I_{1}^{-+})_{0}=i\frac{e^{-i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}\left(\sqrt{\mathfrak{t}}+\frac{\nu}{2\sigma}\right)}\left(1+O\left(\frac{1}{\mathfrak{t}^{\frac{3}{2}}}\right)\right).$ (89)
Expression (89), as well as (87), has no singularities in area (15).
@@
$(I_{1}^{--})_{0}+(I_{1}^{++})_{0}=-\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\varepsilon}\left(1+\delta\right)\\ -\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\varepsilon^{2}}\left[\frac{1}{4\mathfrak{t}}+\frac{1}{2\sqrt{\mathfrak{t}}\,\varepsilon}\left(1+\delta\right)\right].$ (90)
see (52) and (65). Since $\mathfrak{t}$ is a zero of the Airy function’s derivative, $v^{\prime}(\mathfrak{t})=0$, and $\mathfrak{t}\gg 1$, then $\mathfrak{t}^{\frac{3}{2}}=\frac{3}{2}\pi\left(N-\frac{3}{4}\right)+O\left(N^{-1}\right)$, where $N\gg 1$ is the number of the zero $\mathfrak{t}$, see, e.g., [19]. Thus
$\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)=\cos\left(-\frac{\pi}{2}+\pi N+O\left(N^{-1}\right)\right)=O\left(N^{-1}\right)=O\left(\mathfrak{t}^{-\frac{3}{2}}\right).$
At the same time, $\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)=O(1)$. Hence, for different values of $\varepsilon$, the term with $\cos(\ldots)$ can be either of the order of terms with $\sin(\ldots)$, or larger, or smaller. In the area where
$|\varepsilon|\ll\sqrt{\mathfrak{t}}\quad\Leftrightarrow\quad\gamma-\varphi\ll\gamma,$ (91)
i.e. close enough to the limit ray $l_{O}$, the second term with $\sin(\ldots)$ is the main in (90). Not too close to the limit ray, where
$\varepsilon\sim\sqrt{\mathfrak{t}}\quad\Leftrightarrow\quad\gamma-\varphi\sim\gamma,$ (92)
all terms in (90) are of the same order. Above the limit ray, where
$|\varepsilon|\gg\sqrt{\mathfrak{t}}\quad\Leftrightarrow\quad|\gamma-\varphi|\gg\gamma,$ (93)
the term with $\cos(\ldots)$ is the main. Note that conditions (91), (92) and (93) do not contradict inequalities (66) providing smallness of correction terms.
The sum of contributions (87) and (89) of the end point $p=0$ to integrals $I_{1}^{+-}$ and $I_{1}^{-+}$ looks much simpler:
@@
$(I_{1}^{+-})_{0}+(I_{1}^{-+})_{0}=-\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\sqrt{\mathfrak{t}}+\frac{\nu}{2\sigma}}\left(1+O\left(\frac{1}{\mathfrak{t}^{\frac{3}{2}}}\right)\right).$ (94)
It is clear from (50) that the phases of exponentials in (90), (94) agree with the eikonal of cylindrical wave in the area where
$kr\varphi^{4}\ll 1.$ (95)
Now we sum up (90) and (94) and rewrite the resulting expression in polar coordinates (4) using (7) and (53). Matching it with ray formula (2) under conditions (3) and (95), we obtain that the diffracted wave has the form
$u^{\text{dif}}=\left(A_{r}(\varphi;k)+A_{s}(\varphi;k)\right)\frac{e^{ikr}}{\sqrt{kr}},$ (96)
where $A_{r}$ and $A_{s}$ are, respectively, regular and singular parts of the diffraction coefficient. The regular part
@@
$A_{r}(\varphi;k)=-\frac{e^{-i\frac{\pi}{4}}}{\sqrt{2\pi}\mathfrak{t}^{\frac{1}{4}}}\frac{\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\gamma+\varphi}\left(1+O\left(\frac{1}{\mathfrak{t}^{\frac{3}{2}}}\right)\right)$ (97)
has no singularities in the area under consideration. The singular part $A_{s}$ is nonsmooth at the limit ray and is given by following expressions:
@@
$A_{s}(\varphi;k)=-\frac{e^{-i\frac{\pi}{4}}}{2^{\frac{5}{6}}\sqrt{\pi}(ka)^{\frac{2}{3}}\mathfrak{t}^{\frac{3}{4}}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{(\gamma-\varphi)^{3}}\left(1+O\left(\frac{\gamma-\varphi}{\gamma}\right)+\delta\right)$ (98)
under condition (91);
$A_{s}(\varphi;k)=-\frac{e^{-i\frac{\pi}{4}}}{\sqrt{2\pi}\mathfrak{t}^{\frac{1}{4}}}\frac{\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\gamma-\varphi}\left(1+\delta\right)\\ -\frac{e^{-i\frac{\pi}{4}}}{2^{\frac{5}{6}}\sqrt{\pi}\mathfrak{t}^{\frac{1}{4}}}\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)\left[\frac{1}{2\mathfrak{t}(2ka)^{\frac{1}{3}}(\gamma-\varphi)^{2}}+\frac{1}{\sqrt{\mathfrak{t}}\,(ka)^{\frac{2}{3}}(\gamma-\varphi)^{3}}\left(1+\delta\right)\right]$ (99)
under condition (92);
$A_{s}(\varphi;k)=-\frac{e^{-i\frac{\pi}{4}}}{\sqrt{2\pi}\mathfrak{t}^{\frac{1}{4}}}\frac{\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\gamma-\varphi}\left(1+O\left(\frac{\gamma}{\gamma-\varphi}\right)+\delta\right)$ (100)
under condition (93) ($\delta$ is introduced in (65)). Formulas (98), (99) and (100) are applicable in the area where inequalities (66) and (95) hold. Note that the diffraction coefficient is not linear in the jump of curvature $1/a$. This distinguishes the problem under consideration from many problems of diffraction by non-smoothness of curvature (see, e.g., [2, 8, 12]). Such a dependence on curvature jump appeared previously only in the case of $\mathfrak{t}=O(1)$ and the Dirichlet boundary condition [12].

### v2

@@ 5.1.1 Integral $I_{1}^{--}$
$I_{1}^{--}=\frac{1}{4\sqrt{\pi\sigma}}\int_{L_{1}}\frac{e^{if^{--}(p)}}{(\mathfrak{t}-p)^{\frac{1}{4}}}\left(1+\frac{5i}{24(\mathfrak{t}-p)^{\frac{3}{2}}}+O\left(\frac{1}{(\mathfrak{t}-p)^{3}}\right)\right)dp.$
@@
$(I_{1}^{--})_{0}=\frac{e^{-i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\Bigg[\frac{i}{\varepsilon}\left(1+\frac{5i}{24\mathfrak{t}^{\frac{3}{2}}}+O\left(\frac{1}{\mathfrak{t}^{3}}\right)\right)-\frac{1}{4\mathfrak{t}\,\varepsilon^{2}}+\left(\frac{1}{\sigma}-\frac{1}{\sqrt{\mathfrak{t}}}\right)\frac{1}{2\varepsilon^{3}}\left(1+\delta\right)\Bigg].$ (65)
@@
and $\varepsilon$ is introduced in (53). The need to keep so many terms of expansion (65) will be clarified later, when summing the contributions of the end point $p=0$ to the integrals $I_{1}^{\pm\pm}$.
@@
Consider merging of the point $p_{1}$ with the end point, which occurs when $\sigma>\sqrt{\mathfrak{t}}$ and the value of $\varepsilon=\sqrt{\mathfrak{t}}-\nu/2\sigma$ is small, see (53). We expand the phase in a small neighborhood of $p=0$ up to quadratic terms:
@@
Omitting terms of the third order, extending integration from a small neighborhood of $p=0$ to the half-line $0,\infty)$ and passing to the variable
@@
see (53). The value $\zeta_{>}^{2}$ equals, up to the constant summand $2\mathfrak{t}^{3/2}/3$, difference between phase (56) of the wave $u_{2}$ after the rays $\ell_{2}$ pass the caustic and phase (52) of the cylindrical wave. The Fresnel integral in (72) has an imaginary argument, which is atypical of diffraction problems.
@@
The value $\zeta_{<}^{2}$ is equal, up to the constant summand $2\mathfrak{t}^{3/2}/3$, to difference between the phase of the wave $u_{2}$ before the rays $\ell_{2}$ pass the caustic and the phase of the cylindrical wave, see (52) and (56).
@@
holds true. We expand the phase $f^{--}(p)$ in a small neighborhood of $p_{0}$ up to cubic terms:
@@
Omitting terms of the fourth order, extending integration from a small neighborhood of $p_{0}$ to the whole axis and passing to the integration variable $q=(p-p_{0})/2\sigma$, we obtain that the contribution of the critical points of phase to the integral $I_{1}^{--}$ is expressed in terms of the Airy function $v$:
$(I_{1}^{--})_{12}=\frac{e^{-i\left(\frac{2\sigma^{3}}{3}+\sigma(\mathfrak{t}-\nu)-\frac{\widetilde{\nu}^{2}}{4\sigma}\right)}}{2\sqrt{\pi}}\int_{-\infty}^{\infty}e^{i\left(\widetilde{\nu}q-\frac{q^{3}}{3}\right)}\left(1+O\left(\frac{q+q^{4}}{\sigma}\right)\right)dq\\ =e^{-i\left(\frac{2\sigma^{3}}{3}+\sigma(\mathfrak{t}-\nu)\right)}v(-\widetilde{\nu})\left(1+O\left(\frac{1+\widetilde{\nu}^{2}}{\sigma}\right)\right).$ (81)
@@
As above, we expand the phase $f^{--}(p)$ in a neighborhood of $p_{0}=t-\sigma^{2}$ (which is small now) up to cubic terms, see (80). We move terms of the fourth order to the amplitude, extend the integration to the half-line $0,\infty)$ and pass to the variable $q=(p-p_{0})/2\sigma$. Keeping in mind that $\sqrt{\mathfrak{t}}-\sigma$ is small, we come up with
$I_{1}^{--}=-e^{-i\left(\frac{2\sigma^{3}}{3}+\sigma(\mathfrak{t}-\nu)\right)}\mathcal{I}\left(-\widetilde{\nu},\sqrt{\mathfrak{t}}-\sigma\right)\left(1+O\left(\frac{(\sqrt{\mathfrak{t}}-\sigma)^{2}}{\sqrt{\mathfrak{t}}}\right)+O\left(\frac{\widetilde{\nu}^{2}+1}{\sqrt{\mathfrak{t}}}\right)\right).$ (83)
@@ 5.1.2 Integrals $^{++}, I+I^{+-}$ и $I^{-+}$
$(I_{1}^{++})_{0}=i\frac{e^{i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\Bigg[\frac{i}{\varepsilon}\left(1-\frac{5i}{24\mathfrak{t}^{\frac{3}{2}}}+O\left(\frac{1}{\mathfrak{t}^{3}}\right)\right)+\frac{1}{4\mathfrak{t}\,\varepsilon^{2}}+\left(\frac{1}{\sigma}+\frac{1}{\sqrt{\mathfrak{t}}}\right)\frac{1}{2\varepsilon^{3}}\left(1+\delta\right)\Bigg],$ (84)
@@
Comparing the phase of exponential in (85) with expression (50), we conclude that formula (85) describes the wave $u_{1}$, corresponding to the ray family $\ell_{1}$, after reflection from the straight part of the boundary $S_{+}$. Geometrical spreading of these rays is proportional to the square root of the distance from the point $M^{\prime}$ symmetric to the observation point $M$ relative to $S_{+}$, to the caustic $C$, see (38). The approximation (84), (85) is applicable in the area where conditions (67) and
@@
cf. (74) and (76). The quantity $\zeta_{*}^{2}$ is equal, up to the constant summand $2\mathfrak{t}^{3/2}/3$, to the difference between phase (55) of the wave $u_{1}$ after reflection from the straight part of the boundary and phase (52) of the cylindrical wave. It is easy to check that $\sqrt{(x+a\gamma)/a\gamma}=J(M)/J(O)$, where $J(M)$ and $J(O)$ are the values of geometrical spreading of the rays $\ell_{1}$ at the points $M=(x,\gamma x)$ and $O=(0,0)$, respectively, cf. (77).
@@
$(I_{1}^{+-})_{0}=i\frac{e^{i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\Bigg[\frac{i}{\tilde{\varepsilon}}\left(1-\frac{5i}{24\mathfrak{t}^{\frac{3}{2}}}+O\left(\frac{1}{\mathfrak{t}^{3}}\right)\right)+\frac{1}{4\mathfrak{t}\,\tilde{\varepsilon}^{2}}\\ +\left(\frac{1}{\sigma}+\frac{1}{\sqrt{\mathfrak{t}}}\right)\frac{1}{2\tilde{\varepsilon}^{3}}\left(1+O\left(\frac{1}{\mathfrak{t}^{\frac{3}{2}}}\right)+O\left(\frac{1}{\mathfrak{t}\sigma}\right)\right)\Bigg].$ (88)
Here, we introduce the notation
$\tilde{\varepsilon}=\sqrt{\mathfrak{t}}+\frac{\nu}{2\sigma}=2\sqrt{\mathfrak{t}}-\varepsilon.$ (89)
@@
3. The critical point $p_{4}$ of phase of the integral $I^{-+}$ is negative and lies far from the end point $p=0$ for all values of $\sigma$ and $\nu$. Thus, the asymptotics of $I^{-+}$ is given by the end point $p=0$ contribution:
$(I_{1}^{-+})_{0}=\frac{e^{-i\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+i\frac{\nu^{2}}{4\sigma}}}{4\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\Bigg[\frac{i}{\tilde{\varepsilon}}\left(1+\frac{5i}{24\mathfrak{t}^{\frac{3}{2}}}+O\left(\frac{1}{\mathfrak{t}^{3}}\right)\right)-\frac{1}{4\mathfrak{t}\,\tilde{\varepsilon}^{2}}\\ +\left(\frac{1}{\sigma}-\frac{1}{\sqrt{\mathfrak{t}}}\right)\frac{1}{2\tilde{\varepsilon}^{3}}\left(1+O\left(\frac{1}{\mathfrak{t}^{\frac{3}{2}}}\right)+O\left(\frac{1}{\mathfrak{t}\sigma}\right)\right)\Bigg]$ (91)
@@
$(I_{1}^{--})_{0}+(I_{1}^{++})_{0}=-\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\varepsilon}\left(1+\delta\right)\\ -\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\varepsilon}\left[\frac{5}{24\mathfrak{t}^{\frac{3}{2}}}+\frac{1}{4\mathfrak{t}\varepsilon}+\frac{1}{2\sqrt{\mathfrak{t}}\,\varepsilon^{2}}\left(1+\delta\right)\right].$ (92)
see (53) and (66). Deriving the asymptotics for $v^{\prime}(z)$ as $z\to$-$\infty$ from (12), substituting $z=-\mathfrak{t}$ and taking into account that $v^{\prime}(\mathfrak{t})=0$, we obtain following relation:
$\cos\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)=\frac{1}{24\mathfrak{t}^{\frac{3}{2}}}\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)\left(1+O\left(\frac{1}{\mathfrak{t}^{3}}\right)\right)$ (93)
Using (93), we simplify expression (92):
@@
$(I_{1}^{--})_{0}+(I_{1}^{++})_{0}\approx-\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\varepsilon}\left[\frac{1}{4\mathfrak{t}^{\frac{3}{2}}}+\frac{1}{4\mathfrak{t}\varepsilon}+\frac{1}{2\sqrt{\mathfrak{t}}\,\varepsilon^{2}}\right].$ (94)
The sum of end point contributions in $I_{1}^{+-}$ and $I_{1}^{-+}$, see (88) and (91), differs from (94) only by replacing $\varepsilon$ with $\tilde{\varepsilon}$ (89):
@@
$(I_{1}^{+-})_{0}+(I_{1}^{-+})_{0}\approx-\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}\mathfrak{t}^{\frac{1}{4}}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\tilde{\varepsilon}}\left[\frac{1}{4\mathfrak{t}^{\frac{3}{2}}}+\frac{1}{4\mathfrak{t}\tilde{\varepsilon}}+\frac{1}{2\sqrt{\mathfrak{t}}\,\tilde{\varepsilon}^{2}}\right].$ (95)
Finally, summing up (94) and (95), we get
$(I_{1})_{0}\approx-\frac{e^{i\frac{\nu^{2}}{4\sigma}-i\frac{\pi}{4}}}{\sqrt{\pi\sigma}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{\mathfrak{t}^{\frac{1}{4}}}\frac{t+(\nu/2\sigma)^{2}}{(t-(\nu/2\sigma)^{2})^{3}}.$ (96)
It is clear from (51) that the phase of exponential in (96) agree with the eikonal of cylindrical wave in the area where
$kr\varphi^{4}\ll 1,$ (97)
see (51). We rewrite expression (96) in polar coordinates (4) using (7) and (20). Matching it with ray formula (2) under conditions (3) and (97), we obtain that the diffraction coefficient has the form
@@
$A(\varphi;k)\approx 2\sqrt{\frac{2}{\pi}}\frac{\sin\left(\frac{2}{3}\mathfrak{t}^{\frac{3}{2}}+\frac{\pi}{4}\right)}{ka\mathfrak{t}^{\frac{1}{4}}}\frac{\gamma^{2}+\varphi^{2}}{(\varphi^{2}-\gamma^{2})^{3}}e^{-i\frac{\pi}{4}}\approx 2\sqrt{\frac{2}{\pi}}\frac{v(-\mathfrak{t})}{ka}\frac{\gamma^{2}+\varphi^{2}}{(\varphi^{2}-\gamma^{2})^{3}}e^{-i\frac{\pi}{4}},$ (98)
see (12). The expression (98) is applicable in the area where inequality (67) holds and the polar angle is small, $\varphi\ll 1$. The diffraction coefficient (98) is linear in the jump of curvature $1/a$, as it is in other problems of diffraction by jump of curvature (see, e.g., [2, 12, 11, 8]).
In the area where $\varphi\gg\gamma$ the expression (98) becomes
$A(\varphi;k)\approx\sqrt{\frac{2}{\pi}}\frac{v(-\mathfrak{t})}{ka}\frac{2}{\varphi^{4}}e^{-i\frac{\pi}{4}}.$ (99)
The formula (99) coincides with the diffraction coefficient derived in [12], where diffraction of small-number ($\mathfrak{t}=O(1)$) whispering gallery mode was considered.
Moreover, formula (98) agrees with the expression for diffraction coefficient obtained in [9] in the case of non-tangential incidence. Indeed, for a plane incident wave and the Neumann boundary condition the diffraction coefficient $A_{nt}$ has a form (see [9])
$A_{nt}(\varphi,k)=\sqrt{\frac{2}{\pi}}\frac{1}{ka}\frac{\cos\gamma\cos\varphi-1}{(\cos\varphi-\cos\gamma)^{3}}e^{-i\frac{\pi}{4}}.$ (100)
Here, $\gamma$ is the grazing angle of incident wave, $1/a$ is the amplitude of curvature jump and $\varphi$ is the polar angle. Obviously, for small $\gamma$ and $\varphi$ expression (100) matches with (98) up to the factor $v(-\mathfrak{t})/2$.